find the length of the laus rectum of the parabola-


4y^2+2x-20y+17=0


plz help..thanx in advance!!

Grade:

2 Answers

pavani suhani
12 years ago

the above equation can be written as 

                                                   ( 2y-5)^2=2(x+4)

                                                   therefore the length of the latus rectum is 2

Jit Mitra
12 years ago

A small mistake in the above post.

 

(2y-5)2 = -2x-17+25 = -2x+8 = -2(x-4)

 

or, (y- 5/2)2 = -1/2 * (x-4)

 

So Latus Rectum is 1/2 .